aidenj:

A triangle has vertices (-1, 2), (3, 1), and (7, 2). What is the approximate perimeter of the triangle? Round your answer to the nearest hundredth. A. 16 units B. 16.24 units C. 16.31 units D. 17 units

aidenj:

wait if it says round to the nearest hundrenth it gonna have a decimal right

snowflake0531:

1 attachment
snowflake0531:

What is the length of the the top length, count the units

aidenj:

it 16 right

aidenj:

?

snowflake0531:

...

snowflake0531:

If you're just going to guess each of the choices over I"m not playing this with you

aidenj:

im not i counted 16

snowflake0531:

I doubt that

aidenj:

i did bruh since yesterday u bbeen rude u can let it go dude ok

snowflake0531:

Unless you don't know how to count, count the units

snowflake0531:

@aidenj wrote:
i did bruh since yesterday u bbeen rude u can let it go dude ok
I'ts not rude, I just dislike it when you don't actually do it, and just want the answer

aidenj:

i counted 16 like bruh

snowflake0531:

but it's not 16...

aidenj:

if it not 16 what is it

snowflake0531:

1 attachment
aidenj:

my god dude

snowflake0531:

also btw I'm a girl *

aidenj:

now what

snowflake0531:

now we want to find this

1 attachment
aidenj:

it loks like 2.5

snowflake0531:

no, we want to solve it do you remember the pythagorena theorem?

aidenj:

a*2+b*2=c*2

snowflake0531:

It's squared, not times 2

aidenj:

i didnt mean times

aidenj:

so2.5 times 8 right

snowflake0531:

1 attachment
aidenj:

so 2.5 times 2.5 6.25 then what

snowflake0531:

where the literal hell are you getting those numbers

snowflake0531:

First look at a, how long is that

aidenj:

oh 4

snowflake0531:

....bruh, no Look at side a, what is the length

aidenj:

2

snowflake0531:

it's 1 iuhghp4gtihprseyuhiodry6esdr6fti7g8yhu9i

snowflake0531:

then look at B, what is the length of that

aidenj:

4

snowflake0531:

Yes so we can set it up \[4^2+1^2=c^2\] simplify that

aidenj:

4

snowflake0531:

... you could also use a calculator

aidenj:

so 5

snowflake0531:

.... what is 4^2?

aidenj:

8

snowflake0531:

NO, it's not times 2, it's SQUARED

aidenj:

idk

snowflake0531:

4^2 is 4 times itself 2 times, in this case, 4 times 4

aidenj:

16

snowflake0531:

Put this into your calculator: \[8+2\sqrt17\]

aidenj:

16.24

snowflake0531:

there you go

aidenj:

tys